What is the PH of a buffer solution containing 0.2moldm^-3 ethanoic acid and 0.5moldm^-3 sodium ethanoate?Ka=1.75×10^-5 moldm^-3


1
Expert's answer
2021-11-08T08:12:26-0500

.

"PH=-log1.75×10^{-5}+log[\frac{0.2×10^{-3}}{0.5×10^{-3}}]"


"PH=4.757-0.398=4.359"